OpenStudy (anonymous):

Divide 6/13 by 6/12 . A. 12/13 B. 9/16 C. 13/12 D. 1/127

OpenStudy (zzr0ck3r):

\[\dfrac{a}{b}\div \dfrac{c}{d}=\dfrac{a}{b}\times\dfrac{d}{c}=\dfrac{a\times d}{b\times c}\]

OpenStudy (anonymous):

@zzr0ck3r imma be honest with you idk what the hell im looking up lol

OpenStudy (jessie13):

A

Nnesha (nnesha):

look at the example to change division to multiplication multiply top with the `reciprocal` of the bottom fraction